OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2).

OpenStudy (goformit100):

@mary.rojas

HanAkoSolo (jamierox4ev3r):

just plug these two points into the equation \(\huge\frac{y^2-y^1}{x^2-x^1}\)

OpenStudy (anonymous):

These are my choices, not sure how to get to this: y = 5/2 x - 2 y = -5/2 x + 8 Y = 2/5 x - 6 y = -2/5 x + 4

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

just plug in the two points in the equation and that will give you the slope :)

OpenStudy (anonymous):

A right ? @Jamierox4ev3r